# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ 3.1 Backpropagation through graph Laplace equations
$=\left\langle\tau_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
@@
$\tau_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=\delta_{r}(x),$ (29)
@@
$\nabla_{f}\mathcal{J}(r)$ $=\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\tau_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\delta_{r},v\right\rangle_{{\ell^{2}({\mathcal{X}})}}=v(r).$
@@
$\tau_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=0,$ (30)
@@
$=\left\langle h-\delta_{r},\tau_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\nabla_{u}\mathcal{J}(r)$
$=\left\langle h,\tau_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}-\tau_{z}(u(r),r)v(r)$
@@
$=\left\langle\tau_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}-\tau_{z}(u(r),r)v(r)$
@@
$=-\tau_{z}(u(r),r)v(r)-\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)(r)+\nabla_{u}\mathcal{J}(r).$
@@
$\nabla_{\theta}\mathcal{J}=\sum_{x,y\in{\mathcal{X}}}\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}=\sum_{x,y\in{\mathcal{X}}}\frac{1}{2}\left(\nabla_{W}\mathcal{J}(x,y)+\nabla_{W}\mathcal{J}(x,y)\right)\nabla_{\theta}w_{xy}.$ (31)
@@
$\nabla_{\theta}\mathcal{J}=\sum_{x,y\in{\mathcal{X}}}\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}=\sum_{x,y\in{\mathcal{X}}}\frac{1}{2}\left(\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}+\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}\right).$
@@
As in Example 1, we consider the graph $p$-Laplace equation with $p>1$, [^4] The case $p=1$ does not satisfy the continuous differentiability conditions of Theorem 5. $\tau(z,x)=\tau z$, where $\tau$ a nonnegative constant, and $f\in{\ell^{2}({\mathcal{X}})}$. In this case we have $\phi(q,x,y)=|q|^{p-2}q$ and $\phi_{q}(q,x,y)=(p-1)|q|^{p-2}$. The adjoint equation (23) becomes
@@
$\nabla_{g}\mathcal{J}(x)=\nabla_{u}\mathcal{J}(x)-\tau v(x)-(p-1)\operatorname{div}\left(|\nabla u|^{p-2}\nabla v\right)(x)\ \ \text{for}\ x\in\mathcal{L}.$ (34)
When $p=2$, the symmetrized gradient $\nabla_{W}\Psi$ for Laplace learning has the especially simple form $-\nabla u(x,y)\nabla v(x,y)$.
@@ 3.2 Backpropagation through similarity matrices
where $\eta$ is differentiable and $a_{ij}$ is the adjacency matrix of the graph, so $a_{ij}=1$ if there is an edge between nodes $i$ and $j$, and $a_{ij}=0$ otherwise. We consider graphs without self-loops, so $w_{ii}=a_{ii}=0\phantom{=}\forall i$.
@@
$\nabla_{\mathbf{x}_{i}}\mathcal{J}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}v_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})+\sum_{j=1}^{n}b_{j}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})\delta_{ik_{j}}-b_{i}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}}),$ (37)
where $\tilde{g}_{ij}=g_{ij}+g_{ji}$, and the vector $b=(b_{i})_{i}$ and matrix $V=(v_{ij})_{ij}$ are defined by
$b_{i}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}v_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2},\ \ \text{and}\ \ v_{ij}=\eta^{\prime}\left(\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{2\varepsilon_{i}\varepsilon_{j}}\right).$ (38)
@@
$b^{\ell}_{ij}=\frac{v_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})-\frac{v_{ij}}{2\varepsilon_{i}^{2}\varepsilon_{j}^{2}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}\nabla_{\mathbf{x}_{\ell}}(\varepsilon_{i}\varepsilon_{j}).$
@@
$b^{\ell}_{ij}=\frac{v_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})$ $-\frac{v_{ij}}{2\varepsilon_{i}\varepsilon_{j}^{3}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})(\delta_{\ell j}-\delta_{\ell k_{j}})$
$-\frac{v_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})(\delta_{\ell i}-\delta_{\ell k_{i}}).$
@@
$=\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{v_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})$
$\hskip 28.45274pt-\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{v_{ij}}{2\varepsilon_{i}\varepsilon_{j}^{3}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})(\delta_{\ell j}-\delta_{\ell k_{j}})$
$\hskip 28.45274pt-\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{v_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})(\delta_{\ell i}-\delta_{\ell k_{i}})$
$=\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}v_{\ell i}}{\varepsilon_{\ell}\varepsilon_{i}}(\mathbf{x}_{\ell}-\mathbf{x}_{i})-\left[\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}v_{\ell i}}{2\varepsilon_{\ell}^{3}\varepsilon_{i}}\|\mathbf{x}_{\ell}-\mathbf{x}_{i}\|^{2}\right](\mathbf{x}_{\ell}-\mathbf{x}_{k_{\ell}})$
$\hskip 180.67499pt-\sum_{i=1}^{n}\left[\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}v_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}\right](\mathbf{x}_{k_{i}}-\mathbf{x}_{i})\delta_{\ell k_{i}},$
@@
$\nabla_{\mathbf{x}_{\ell}}\mathcal{J}=\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}v_{\ell i}}{\varepsilon_{\ell}\varepsilon_{i}}(\mathbf{x}_{\ell}-\mathbf{x}_{i})+\sum_{i=1}^{n}b_{i}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})\delta_{\ell k_{i}}-b_{\ell}(\mathbf{x}_{\ell}-\mathbf{x}_{k_{\ell}}).$ (40)
@@
$\nabla_{\mathbf{x}_{i}}\mathcal{J}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}v_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})+\sum_{j=1}^{n}c_{ij}b_{j}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}}).$ (41)
@@ 3.2.1 Sparse and efficient gradient computations
We use the expression (41) from Remark 10. The matrices $A$ (adjacency matrix), $V$ (from Lemma 8), and $C$ (from Remark 10) require minimal additional computational overhead, as their structure and values all come from the $k$-NN search and weight matrix construction required in the forward pass to construct the weight matrix $W$. In our experiments, for example, we use $\eta(z)=\exp(-4z)$, so $V=-4W$. The matrix $G$ is computed as described in Section 3.1.
@@
$h_{ij}=\frac{\tilde{g}_{ij}a_{ij}v_{ij}}{\varepsilon_{i}\varepsilon_{j}}$

### v2

@@ 3.1.2 Nonlinear elliptic equations on graphs
$=\left\langle\sigma_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}$
@@
$\sigma_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=\delta_{r}(x),$ (36)
@@
$\nabla_{f}\mathcal{J}(r)$ $=\left\langle h,\nabla_{u}\mathcal{J}\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\sigma_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}=\left\langle\delta_{r},v\right\rangle_{{\ell^{2}({\mathcal{X}})}}=v(r).$
@@
$\sigma_{z}(u(x),x)h(x)+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right)=0,$ (37)
@@
$=\left\langle h-\delta_{r},\sigma_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}+\nabla_{u}\mathcal{J}(r)$
$=\left\langle h,\sigma_{z}(u,\cdot)v+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)\right\rangle_{{\ell^{2}({\mathcal{X}})}}-\sigma_{z}(u(r),r)v(r)$
@@
$=\left\langle\sigma_{z}(u,\cdot)h+\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla h\right),v\right\rangle_{{\ell^{2}({\mathcal{X}})}}-\sigma_{z}(u(r),r)v(r)$
@@
$=-\sigma_{z}(u(r),r)v(r)-\operatorname{div}\left(\phi_{q}(\nabla u,\cdot)\nabla v\right)(r)+\nabla_{u}\mathcal{J}(r).$
We finally use that $v(r)=0$ for $r\in\mathcal{L}$.
@@
$\nabla_{\theta}\mathcal{J}=\sum_{x,y\in{\mathcal{X}}}\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}=\sum_{x,y\in{\mathcal{X}}}\frac{1}{2}\left(\nabla_{W}\mathcal{J}(x,y)+\nabla_{W}\mathcal{J}(y,x)\right)\nabla_{\theta}w_{xy}.$ (38)
@@
$\nabla_{\theta}\mathcal{J}=\sum_{x,y\in{\mathcal{X}}}\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}=\sum_{x,y\in{\mathcal{X}}}\frac{1}{2}\left(\nabla_{W}\mathcal{J}(x,y)\nabla_{\theta}w_{xy}+\nabla_{W}\mathcal{J}(y,x)\nabla_{\theta}w_{yx}\right).$
@@
As in Example 1, we consider the graph $p$-Laplace equation with p>1p>1, [^4] The case $p=1$ does not satisfy the continuous differentiability conditions of Theorem 5. $\sigma(z,x)=\tau z$, where $\tau$ a nonnegative constant, and $f\in{\ell^{2}({\mathcal{X}})}$. In this case we have $\phi(q,x,y)=|q|^{p-2}q$ and $\phi_{q}(q,x,y)=(p-1)|q|^{p-2}$. The adjoint equation (28) becomes
@@
$\nabla_{g}\mathcal{J}(x)=\nabla_{u}\mathcal{J}(x)-(p-1)\operatorname{div}\left(|\nabla u|^{p-2}\nabla v\right)(x)\ \ \text{for}\ x\in\mathcal{L}.$ (41)
When $p=2$, the symmetrized gradient $\nabla_{W}\Psi$ for Laplace learning has the especially simple form $\nabla u(x,y)\nabla v(x,y)$.
@@ 3.2 Backpropagation through similarity matrices
where $\eta$ is differentiable and $a_{ij}$ is the adjacency matrix of the graph, so $a_{ij}=1$ if there is an edge between nodes $i$ and $j$, and $a_{ij}=0$ otherwise. We consider graphs without self-loops, so $w_{ii}=a_{ii}=0$ for all $i$.
@@
$\nabla_{\mathbf{x}_{i}}\mathcal{J}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}p_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})+\sum_{j=1}^{n}b_{j}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})\delta_{ik_{j}}-b_{i}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}}),$ (44)
where $\tilde{g}_{ij}=g_{ij}+g_{ji}$, and the vector $b=(b_{i})_{i}$ and matrix $P=(p_{ij})_{ij}$ are defined by
$b_{i}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}p_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2},\ \ \text{and}\ \ p_{ij}=\eta^{\prime}\left(\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{2\varepsilon_{i}\varepsilon_{j}}\right),$ (45)
where $\eta^{\prime}$ is the derivative of $\eta$.
@@
$b^{\ell}_{ij}=\frac{p_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})-\frac{p_{ij}}{2\varepsilon_{i}^{2}\varepsilon_{j}^{2}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}\nabla_{\mathbf{x}_{\ell}}(\varepsilon_{i}\varepsilon_{j}).$
@@
$b^{\ell}_{ij}=\frac{p_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})$ $-\frac{p_{ij}}{2\varepsilon_{i}\varepsilon_{j}^{3}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})(\delta_{\ell j}-\delta_{\ell k_{j}})$
$-\frac{p_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})(\delta_{\ell i}-\delta_{\ell k_{i}}).$
@@
$=\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{p_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})(\delta_{i\ell}-\delta_{j\ell})$
$\hskip 28.45274pt-\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{p_{ij}}{2\varepsilon_{i}\varepsilon_{j}^{3}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}})(\delta_{\ell j}-\delta_{\ell k_{j}})$
$\hskip 28.45274pt-\sum_{i,j=1}^{n}g_{ij}a_{ij}\frac{p_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})(\delta_{\ell i}-\delta_{\ell k_{i}})$
$=\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}p_{\ell i}}{\varepsilon_{\ell}\varepsilon_{i}}(\mathbf{x}_{\ell}-\mathbf{x}_{i})-\left[\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}p_{\ell i}}{2\varepsilon_{\ell}^{3}\varepsilon_{i}}\|\mathbf{x}_{\ell}-\mathbf{x}_{i}\|^{2}\right](\mathbf{x}_{\ell}-\mathbf{x}_{k_{\ell}})$
$\hskip 180.67499pt-\sum_{i=1}^{n}\left[\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}p_{ij}}{2\varepsilon_{i}^{3}\varepsilon_{j}}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}\right](\mathbf{x}_{k_{i}}-\mathbf{x}_{i})\delta_{\ell k_{i}},$
@@
$\nabla_{\mathbf{x}_{\ell}}\mathcal{J}=\sum_{i=1}^{n}\frac{\tilde{g}_{\ell i}a_{\ell i}p_{\ell i}}{\varepsilon_{\ell}\varepsilon_{i}}(\mathbf{x}_{\ell}-\mathbf{x}_{i})+\sum_{i=1}^{n}b_{i}(\mathbf{x}_{i}-\mathbf{x}_{k_{i}})\delta_{\ell k_{i}}-b_{\ell}(\mathbf{x}_{\ell}-\mathbf{x}_{k_{\ell}}).$ (47)
@@
$\nabla_{\mathbf{x}_{i}}\mathcal{J}=\sum_{j=1}^{n}\frac{\tilde{g}_{ij}a_{ij}p_{ij}}{\varepsilon_{i}\varepsilon_{j}}(\mathbf{x}_{i}-\mathbf{x}_{j})+\sum_{j=1}^{n}c_{ij}b_{j}(\mathbf{x}_{j}-\mathbf{x}_{k_{j}}).$ (48)
@@ 3.2.1 Sparse and efficient gradient computations
We use the expression (48) from Remark 10. The matrices $A$ (adjacency matrix), $P$ (from Lemma 8), and $C$ (from Remark 10) require minimal additional computational overhead, as their structure and values all come from the $k$-NN search and weight matrix construction required in the forward pass to construct the weight matrix $W$. In our experiments, for example, we use $\eta(z)=\exp(-4z)$, so P=4WP=-4W. The matrix $G$ is computed as described in Section 3.1.
@@
$h_{ij}=\frac{\tilde{g}_{ij}a_{ij}p_{ij}}{\varepsilon_{i}\varepsilon_{j}}$